# Assessing the Feasibility of Internet-Sourced Video for Automatic Cattle Lameness Detection

### v2

@@ Abstract
Cattle lameness is often caused by hoof injuries or interdigital dermatitis, leads to pain and significantly impacts essential physiological activities such as walking, feeding, and drinking. This study presents a deep learning-based model for detecting cattle lameness, sickness, or gait abnormalities using publicly available video data. The dataset consists of 50 unique videos from 42 individual cattle, recorded from various angles in both indoor and outdoor environments. Half of the dataset represents naturally walking (normal/non-lame) cattle, while the other half consists of cattle exhibiting gait abnormalities (lame). To enhance model robustness and generalizability, data augmentation was applied to the training data. The pre-processed videos were then classified using two modified deep learning models: ConvLSTM2D and 3D CNN. A comparative analysis of the results demonstrates strong classification performance. Specifically, the 3D CNN model achieved a video-level classification accuracy of 90%, with precision, recall, and f1-score of 90.9%, 90.9%, and 90.91% respectively. The ConvLSTM2D model exhibited a slightly lower accuracy of 85%. This study highlights the effectiveness of directly applying classification models to learn spatiotemporal features from video data, offering an alternative to traditional multi-stage approaches that typically involve object detection, pose estimation, and feature extraction. Besides, the findings demonstrate that the proposed deep learning models, particularly the 3D CNN, effectively classify and detect lameness in cattle while simplifying the processing pipeline.
Index Terms:
Cattle, Lameness detection, Deep-learning, Computer vision techniques, Image processing.
I Introduction
Figure 1: Overall workflow of the cattle lameness classification model
Cattle farms significantly contribute to global food security by providing meat and dairy products, with continuously increasing demand. While large-scale farming is now common, it presents operational challenges and necessitates continuous monitoring of animal health and welfare, as this directly impacts farm productivity [1, 2]. However, one of the leading health issues in cattle is lameness, a common and painful disorder frequently caused by gait-related injuries. Research indicates that approximately one in four cattle is affected by lameness at any given time, which reduces farm productivity and plays a substantial role in increasing meat and dairy prices [3, 5, 4]. A common practice about identifying lame cattle is herd locomotion scoring which requires visual observation of each animal individually [6, 7, 8, 9]. This process is cost and time consuming since they tend to increase human commissioning. Therefore, it is highly necessary to deploy advanced automatic technologies for identification of such health difficulties.
In recent years, AI through ML and data analysis have come up with smart solutions to serve large-scale cattle farming in different ways, e.g., by early detection of diseases and infections, by improving the production characteristics, etc [10]. AI uses data collected from livestock farms using various advanced tools including cameras, drones, robotics, and relevant IoT devices [14]. Although manual video cameras have been commonly used for a long time to surveillance the activities of animals from farmers’ desks [11, 12], nowadays, computer vision and pattern recognition techniques have added new opportunities using the video data to automatically monitor animals’ health and welfare 24/7 [13]. Whether using a manual or computer vision (CV) approach, lameness is detected by observing behavioral changes in cattle—how they react and respond during regular activities in their living environment. Advanced deep learning-oriented AI models analyze image or video data from different perspectives, depending on the animal’s behavior (e.g., eating, walking, and general movement) [15, 16]. Abnormal behavior is often readily apparent when cattle walk, as lameness is closely associated with painful gait and foot disease. The automation primarily focuses on detecting deviations between normal and abnormal behavior, classifying them to make decisions and further analysis.
However, the impact of such development in livestock farming is enormous, it reduces the labor costs, helps to identify health issues immediately, and increases production efficiency. Therefore, researchers are continuously working on the innovative smart farming solution with image processing and CV.
Fuentes et al. [14] used two Hikvision surveillance cameras to collect video data for their study. The cameras captured video from two different angles and extracted frames based on animal activity. They then used these frames to create a dataset, which was employed to train a model that effectively monitored the cows and performed statistical analysis to capture behavioral changes. McDonagh et al. [8] introduced a neural network model for behavior prediction by monitoring 46 individual pregnant cows. They specially considered the 10 hours prior video data to calving and focused on seven different behaviors. The model showed successful classification of the behavior over 80% instances and was able to predict calving before several hours with 83% accuracy. Fernández-Carrión et al. [10] proposed a deep learning model for the early detection of disease in Eurasian wild boar by monitoring their motion. They found a negative correlation between the animals’ movement and fever caused by the disease; the movement of infected animals was significantly reduced compared to that of healthy ones. In another study, Qiao et al. [15] proposed a combined model of Convolutional 3D and Convolutional Long Short-Term Memory (LSTM), which achieved higher accuracy compared to other state-of-the-art models in classifying five behaviors. Maria et al. [17]. focused on swine health, specifically identifying respiratory diseases. They considered eye temperature, heart rate, and respiration rate by extracting information from thermal infrared images. Their results demonstrated the successful implementation of CV in the early detection of respiratory diseases by considering physiological changes in pigs.
In another article, Barney et al. [19]. proposed a modified Mask-RCNN model for real-time automated lameness detection with high accuracy. In [14], the authors used Mask-RCNN and YOLO, both object detection algorithms. After detecting the cattle, they assigned a local ID to track each animal, then extracted features, and applied machine learning (ML) algorithms. Similarly, Russello et al. [5] presented a deep learning model to detect lameness by extracting nine annotations from video data. Wu et al. [18] introduced an LSTM model for large-scale farming to identify lame cows automatically. They used 50 videos, with a lame-to-non-lame sample ratio of 20:30. From these videos, they randomly selected 5,000 frames for their experiment, focusing solely on the position and movement of the cows’ legs. Their LSTM-based classification model achieved 98.57% accuracy, outperforming three traditional machine learning algorithms using a 50:50 train-test split.
However, the literature shows that cattle lameness was investigated in previous studies using surveillance video data collected from farms, with cameras installed at various angles. Additionally, most of those studies focused on small samples, and the investigations were conducted within limited and specific contexts. This study aimed to investigate lameness detection in large cattle farms, where hundreds of cattle need to be managed simultaneously. The research was conducted in two parts: video data collection and the development of a classification model. To achieve this, YouTube videos containing both lame and non-lame cattle were used. The labeling of normal and lame cattle was performed based on the title and description of each video. Since online videos were captured from diverse environments and locations, dataset diversity and robustness were ensured. To enhance the dataset size, vertical augmentation was applied to the training data. Moreover, to build the classification model, two deep learning classifiers, ConvLSTM2D and 3D CNN were utilized. The models were evaluated using multiple performance metrics. The classification results indicate successful implementation of online data with the proposed models, achieving comparable or superior accuracy compared to state-of-the-art models.
@@ I Introduction
The remainder of this paper is organized as follows: Section II presents the materials and methods, including data collection, preprocessing, and model construction. Section III discusses the results obtained from the proposed models. Section IV discusses the findings and compares them with existing studies, along with limitations and future directions. Finally, Section V concludes the paper.
II Materials and Methods
@@
II-A Dataset
@@ II-A1 Data collection
II-A2 Basic Information
TABLE I: Summary of the dataset statistics.
@@ II-A2 Basic Information
To create the dataset, a total of 50 videos featuring 42 cattle were considered, with an average duration of 5.67 seconds per video. We ensured an equal class distribution of the data, meaning that 25 lame and 25 nomal cattle videos were included. Table I presents the dataset distribution, showing the number of videos, total duration in seconds, and the corresponding frame counts for the normal and lame classes. The normal class videos have a cumulative duration of 134.46 seconds and consist of 4050 frames, whereas the lame class videos have a total duration of 148.91 seconds and comprise 5900 frames. Overall, the dataset includes 50 videos with a combined duration of 283.37 seconds and a total of 9950 frames. Besides, Figure LABEL:fig:all_images displays example frames from cattle video sequences, with upper three images representing lame cattle and the other representing normal cattle. The original videos had varying resolutions due to collection from different online sources. To standardize the input data, all videos were resized to a resolution of $500\times 500$ pixels. The dataset used in this study is available at the following Google Drive link: https://drive.google.com/drive/folders/1JHutEYjR6tAvXNnJuMhVgJabNC8tq7z7?usp=sharing.
@@
0:
  • $\bullet$
@@
  • $\bullet$
    $N_{\text{frames}}$: Number of frames per video
  • $\bullet$
    $S_{\text{target}}$: Target size for frame resizing
  • $\bullet$
    $p_{\text{aug}}$: Data augmentation probability
  • $\bullet$
    Batch size $=B$, Number of epochs $=E$
0:  Trained 3DCNN or ConvLSTM2D model; performance metrics
1:  1. Data Preparation
2:  for $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{test}}$ do
@@
4:     if $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}$ then
5:        $(X,y)\leftarrow(X\cup\{F_{i}\},y\cup\{l_{i}\})$
6:     else
7:        $(X_{\text{test}},y_{\text{test}})\leftarrow(X_{\text{test}}\cup\{F_{i}\},y_{\text{test}}\cup\{l_{i}\})$
8:     end if
9:  end for
10:  Apply horizontal flip augmentation to $X_{\text{train}}$ with probability $p_{\text{aug}}$
11:  Normalize $X_{\text{train}},X_{\text{test}}$ and one-hot encode labels
12:  2. Model Training
13:  $M\leftarrow\text{CreateModel}(N_{\text{frames}},S_{\text{target}})$
14:  $M\leftarrow\text{TrainModel}(M,X_{\text{train}},y_{\text{train}},B,E)$
15:  3. Model Evaluation
16:  $\hat{y}\leftarrow M(X_{\text{test}})$
17:  metrics,confusion $\text{metrics},\text{confusion matrix}\leftarrow\text{Evaluate}(y_{\text{test}},\hat{y})$
18:  return  $M$, metrics, confusion matrix
II-B Data used in this study
TABLE II: Basic information of the dataset used in this study compared with the most relevant datasets in the literature
Study Train Frames Test Frames Test Environment Focused Activity
1 1376 345 Within the barn Intensive farming
2 1000 500 Passing alley Intensive farming
3 490 210 Walked along a narrow path Intensive farming
4 690 341 Barn to milking parlor pathway Intensive farming
5 600 120 Corridor after the milking parlor Intensive farming
Table II provides an overview of the data used in this research alongside datasets from previous cattle lameness detection studies. The table compares these studies based on the number of training and testing frames, the test environment, and the focused activity. As previously mentioned, the dataset consists of a total of 50 videos, including 25 normal and 25 lame cattle videos. Out of these, 30 videos (15 normal and 15 lame) were used for training the model, while the remaining 20 videos (10 normal and 10 lame) were used for testing. Previous studies focused on intensive farming environments, where cattle were primarily managed in indoor grazing conditions and data was collected from specific, limited locations such as within a barn, a passing alley, a narrow path, and corridors related to milking parlors. In contrast, this study focuses on a semi-intensive farming setting and incorporates both indoor and outdoor environments. This approach captures a wider range of cattle movement behaviors, lighting conditions, and surface variations, potentially leading to a more generalizable model for cattle lameness detection.
II-B1 Data pre-processing
Preprocessing in this investigation was conducted in several distinct steps, including frame extraction and data augmentation. The overall average frame rate was approximately 35.12 frames per second. However, due to the high volume of frames, only 25 images were randomly selected from each training video, resulting in a total of 750 images. These images were then cropped to a 224×224 pixel ratio. However, data augmentation is a widely recognized method for addressing limited data issues and helps to create robust and sustainable classification models. Numerous data augmentation techniques are available, as demonstrated by Jia et al. [3], who employed nine different methods, such as Mixup, Cutout, CutMix, etc. This study considered only image flipping as an augmentation technique, applied solely to the training data. All 30 training videos (750 frames) were virtually flipped, generating a total of 1500 images. In test data, 500 images were extracted from 20 videos. However, no data augmentation was performed on the test dataset to preserve integrity and ensure a fair assessment of the model’s generalization ability.
II-C Model Construction
@@
The algorithm 1 outlines the pseudocode of the classfication model. It begins with data preparation by processing the training and testing datasets ($\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{test}}$), extracting frames from each video, resizing them to a target size ($S_{\text{target}}$), and ensuring each video has $N_{\text{frames}}$ frames through padding or truncation, creating input data ($X$) and labels ($y$). Training data is augmented with random horizontal flipping (probability $p_{\text{aug}}$). Pixel values are normalized, and behavior labels (“normal” and “lame”) are encoded numerically. Next, model training involves creating a 3D CNN or ConvLSTM2D model, configured using $N_{\text{frames}}$ and $S_{\text{target}}$, and training it on the prepared data ($X_{\text{train}}$, $y_{\text{train}}$) using a batch size ($B$) and epochs ($E$). The model learns patterns indicative of different behavior classes. Finally, model evaluation uses the trained 3D CNN model to predict behavior labels ($\hat{y}$) for the testing dataset ($X_{\text{test}}$), comparing them to true labels ($y_{\text{test}}$) to assess performance. Metrics like accuracy are calculated, and a confusion matrix is generated. The trained model, performance metrics, and confusion matrix are returned as output.
II-C1 3D CNN

### v3

# Direct Video-Based Spatiotemporal Deep Learning for Cattle Lameness Detection

@@ Abstract
Cattle lameness is a prevalent health problem in livestock farming, often resulting from hoof injuries or infections, and severely impacts animal welfare and productivity. Early and accurate detection is critical for minimizing economic losses and ensuring proper treatment. This study proposes a spatiotemporal deep learning framework for automated cattle lameness detection using publicly available video data. We curate and publicly release a balanced set of 50 online video clips featuring 42 individual cattle, recorded from multiple viewpoints in both indoor and outdoor environments. The videos were categorized into lame and non-lame classes based on visual gait characteristics and metadata descriptions. After applying data augmentation techniques to enhance generalization, two deep learning architectures were trained and evaluated: 3D Convolutional Neural Networks (3D CNN) and Convolutional Long-Short-Term Memory (ConvLSTM2D). The 3D CNN achieved a video-level classification accuracy of 90%, with a precision, recall, and F1 score of 90. 9% each, outperforming the ConvLSTM2D model, which achieved 85% accuracy. Unlike conventional approaches that rely on multistage pipelines involving object detection and pose estimation, this study demonstrates the effectiveness of a direct end-to-end video classification approach. Compared with the best end-to-end prior method (C3D-ConvLSTM, 90. 3%), our model achieves comparable accuracy while eliminating pose estimation pre-processing.The results indicate that deep learning models can successfully extract and learn spatio-temporal features from various video sources, enabling scalable and efficient cattle lameness detection in real-world farm settings.
keywords:
keyword1, Keyword2, Keyword3, Keyword4
1 Introduction
Livestock production is a cornerstone of global food security, and the dairy and beef sectors depend critically on healthy, mobile animals. Lameness—typically arising from hoof lesions, interdigital dermatitis, or gait-related injuries—remains one of the most prevalent welfare and economic challenges in cattle farming. A recent systematic review by [1] estimates that 23–25% of dairy cows worldwide exhibit clinical or sub-clinical lameness at any given time, leading to reduced milk yield, delayed reproduction, and increased culling rates. Because lameness pain alters key behaviours such as walking, standing, and feeding, early detection is essential for prompt treatment and cost containment.
Traditional detection relies on locomotion scoring, where trained personnel visually assess each animal. Although effective for small herds, this method is labour-intensive, subjective, and often infeasible on large commercial farms. Recent advances in artificial intelligence (AI) and computer vision (CV) have enabled a suite of automated solutions for animal-health monitoring. Modern farms now deploy surveillance cameras, drones, and IoT sensors that continuously capture behavioural data [15, 14]. Deep learning, in particular, can extract spatiotemporal patterns from video streams, facilitating early diagnosis of lameness, calving, mastitis, and other conditions [16, 17].
Prior studies demonstrate that convolutional neural networks (CNNs), key-point detectors, and pose-estimation pipelines can achieve high frame-level lameness classification accuracy in controlled settings [20, 6]. These systems reduce subjective bias, operate continuously, and potentially lower veterinary costs.
Nevertheless, three limitations persist:
  • Limited environmental diversity. Most datasets are captured in single barns or passage alleys under homogeneous lighting and flooring conditions, hindering generalisation to other farms.
  • Multi-stage complexity. Many pipelines require sequential object detection, tracking, pose estimation, and handcrafted feature extraction, increasing latency and engineering burden.
  • Restricted public availability. Few lameness video datasets are openly released, which slows reproducibility and benchmarking.
This study addresses these gaps by investigating whether a direct, end-to-end spatiotemporal deep-learning model can learn discriminatory gait patterns from freely available internet videos filmed under diverse conditions. Our key contributions include:
  1. 1.
    Open and diverse dataset. We curate and publicly release a balanced set of 50 internet video clips containing 42 individual cattle, captured from multiple viewpoints in both indoor and outdoor scenes.
  2. 2.
    End-to-end deep learning architecture. We design and compare two spatiotemporal models—3D Convolutional Neural Networks (3D-CNN) and Convolutional Long Short-Term Memory networks (ConvLSTM2D)—that operate directly on video tensors without intermediate object-detection or pose-estimation steps.
  3. 3.
    Comprehensive evaluation. Using a strict train–test split, the proposed 3D-CNN attains 90% video-level accuracy, matching or surpassing state-of-the-art multi-stage pipelines while reducing system complexity. We provide the dataset link, annotated labels, and source code to foster transparent benchmarking and future research.
Unlike traditional CNNs that analyze static spatial features, 3D Convolutional Neural Networks extend their learning across both spatial and temporal axes, allowing the model to directly capture motion-based cues from video sequences. This property is especially beneficial for detecting lameness, where subtle irregularities in gait unfold over time. By leveraging 3D convolutions, our model can learn temporal patterns of movement without relying on handcrafted features or frame-wise annotations, enabling more accurate and scalable cattle monitoring in real-world environments.
@@ 1 Introduction
The remainder of this paper is organized as follows. The next section 2 surveys the related literature. Section 3 presents the materials and methods, including data collection, preprocessing, and model construction. Section 4 discusses the results obtained from the proposed models. Section 5 discusses the findings and compares them with existing studies, along with limitations and future directions. Finally, Section 6 concludes the paper.
2 Related Work
Cattle farming plays a crucial role in ensuring global food security by supplying essential meat and dairy products. With rising global demand, large-scale cattle farming has become increasingly common. However, this trend introduces operational challenges, particularly in monitoring animal health and welfare, which are directly tied to productivity outcomes [2, 3, 25]. One of the most prevalent and costly health issues in cattle is lameness—a painful condition often linked to gait-related injuries or hoof disorders. Studies estimate that approximately one in four cattle are affected by lameness at any given time, resulting in reduced productivity and higher production costs, which can contribute to elevated market prices for meat and dairy [4, 6, 5].
A widely used approach to detect lameness is herd locomotion scoring, which involves the visual assessment of each animal [7, 8, 9, 10]. While this method is informative, it is labor-intensive, time-consuming, and subject to human bias. These limitations highlight the need for automated, scalable technologies capable of early and accurate identification of lameness in large herds.
In recent years, artificial intelligence (AI), particularly machine learning (ML) and data-driven techniques, has introduced promising solutions for automating animal health monitoring. These technologies support large-scale farming by enabling early disease detection, optimizing animal welfare, and enhancing productivity [11]. Data are typically gathered using advanced tools such as fixed-position cameras, drones, mobile robots, and Internet of Things (IoT) sensors [15]. While traditional surveillance cameras have long been used by farmers to passively monitor livestock [12, 13], the integration of computer vision (CV) and pattern recognition algorithms has enabled continuous, automated monitoring of livestock health and behavior [14].
Both manual and automated CV-based approaches rely on observing behavioral cues to identify lameness. Lameness typically manifests as altered locomotion, reduced activity, or abnormal postures—behaviors that can be detected during routine activities like walking, standing, or feeding. Modern deep learning models can analyze video data from multiple perspectives to capture such behavioral cues and distinguish between normal and abnormal gait patterns [16, 17]. These models focus on learning the spatial and temporal dynamics of movement, allowing automated systems to detect early deviations and classify them accordingly.
The growing application of such models in livestock monitoring has significantly reduced reliance on manual inspection, lowered labor costs, and increased both detection speed and accuracy. Consequently, research in this area continues to advance, with increasing efforts to develop robust, scalable, and data-efficient AI-based farming solutions.
For instance, Fuentes et al. [15] employed two Hikvision surveillance cameras to record cattle activity from different angles. They extracted frames from the videos to build a dataset, which was used to train a classification model capable of monitoring cow behavior and detecting changes. McDonagh et al. [9] developed a neural network to monitor 46 pregnant cows, analyzing ten hours of pre-calving video per subject. Their model successfully identified seven distinct behaviors and achieved over 80% classification accuracy, predicting calving with 83% precision several hours in advance.
Similarly, Fernández-Carrión et al. [11] monitored the motion of Eurasian wild boar to detect early signs of disease. Their model demonstrated a significant inverse correlation between reduced movement and the onset of fever, suggesting the feasibility of motion-based disease detection. Qiao et al. [16] introduced a hybrid model combining 3D CNN and ConvLSTM layers, outperforming state-of-the-art baselines in classifying five cow behaviors. In the swine domain, Maria et al. [18] applied thermal imaging and CV to detect respiratory illnesses by analyzing physiological features such as heart rate, eye temperature, and breathing patterns.
Additional approaches have explored object detection and pose estimation pipelines. Barney et al. [20] proposed a modified Mask R-CNN architecture for real-time lameness detection using multi-cattle pose estimation. In a similar context, researchers in [19] combined object detection models (Mask R-CNN, YOLO) with post-hoc machine learning classifiers. Each detected cow was assigned a local identifier for tracking, followed by feature extraction and classification. Russello et al. [6] designed a deep learning model that extracted nine locomotion traits from videos and used them to detect lameness. In their study, Wu et al. [19] used 50 cattle videos—20 lame and 30 non-lame—and randomly sampled 5,000 frames, focusing on leg movement. Their LSTM model achieved a classification accuracy of 98.57%, outperforming traditional ML baselines under a 50:50 train–test split.
Despite promising results, many of these studies relied on data captured in restricted, controlled farm environments, often using small sample sizes and homogeneous conditions. The generalizability of such models to real-world settings, where lighting, surface conditions, and animal orientation vary, remains limited. Furthermore, most existing pipelines involve multiple processing stages—such as object detection, tracking, and pose estimation—which increase complexity and computational overhead.
In response to these limitations, the current study explores an alternative approach: direct end-to-end classification using unstructured online video data. We collected 50 publicly available YouTube videos featuring 42 individual cattle with diverse visual characteristics. Each video was labeled as either lame or non-lame based on metadata and gait observation. Since the videos were recorded under varying environmental conditions and viewpoints, the dataset offers enhanced variability and robustness. To augment the training set, horizontal flipping was applied to increase data diversity.
We then trained two spatiotemporal deep learning models—3D CNN and ConvLSTM2D—on this dataset and evaluated their performance using standard classification metrics. The results demonstrate that our approach, which avoids multi-step feature engineering, can effectively classify cattle lameness with performance comparable to or exceeding that of existing state-of-the-art models. Our findings highlight the potential of internet-sourced video data and simplified model pipelines for scalable livestock health monitoring in real-world scenarios.
Table 1 shows a comparison table summarizing key related works on cattle lameness detection, highlighting aspects such as data source, method type, model, sample size, and reported accuracy.
Table 1: Comparison of Related Work on Cattle Lameness Detection
Ref. Data Source Model Type Sample Size Accuracy (%)
[5] 3D cameras in barns Logistic Regression with extracted features $\sim$1700 frames 81.2
[19] Barn videos, object detection via YOLO LSTM with leg motion features 5000 frames 98.6
[16] Controlled video environment C3D + ConvLSTM hybrid model 15 cows 90.3
[20] Multi-cattle surveillance video Modified Mask-RCNN for pose tracking Not reported
[6] Farm video data Pose estimation + locomotion traits 80.1
[9] Pre-calving cow monitoring CNN-based behavior classifier 46 cows 83.0
This study YouTube videos (diverse environments) End-to-end 3D CNN / ConvLSTM2D 50 videos 90.0 / 85.0
3 Materials and Methods
@@
Figure 1: Overall workflow of the cattle lameness classification model.
3.1 Dataset
@@ 3.1.1 Data collection
Figure 2: Example frames collected from video sequences of cattle used in this study. The upper set depicts frames from lame cattle, whereas the lower set illustrates frames from normal cattle.
3.1.2 Basic Information
Table 2: Summary of the dataset statistics.
@@ 3.1.2 Basic Information
To create the dataset, a total of 50 videos featuring 42 cattle were considered, with an average duration of 5.67 seconds per video. We ensured an equal class distribution of the data, meaning that 25 lame and 25 nomal cattle videos were included. Table 2 presents the dataset distribution, showing the number of videos, total duration in seconds, and the corresponding frame counts for the normal and lame classes. The normal class videos have a cumulative duration of 134.46 seconds and consist of 4050 frames, whereas the lame class videos have a total duration of 148.91 seconds and comprise 5900 frames. Overall, the dataset includes 50 videos with a combined duration of 283.37 seconds and a total of 9950 frames. Besides, Figure 2 displays example frames from cattle video sequences, with upper three images representing lame cattle and the other representing normal cattle. The original videos had varying resolutions due to collection from different online sources. To standardize the input data, all videos were resized to a resolution of $500\times 500$ pixels.
3.2 Data used in this study
Table 3: Basic information of the dataset used in this study compared with the most relevant datasets in the literature
Study Train Frames Test Frames Test Environment Focused Activity
[26] 1376 345 Within the barn Intensive farming
[10] 1000 500 Passing alley Intensive farming
[19] 490 210 Walked along a narrow path Intensive farming
[21] 690 341 Barn to milking parlor pathway Intensive farming
[4] 600 120 Corridor after the milking parlor Intensive farming
Table 3 provides an overview of the data used in this research alongside datasets from previous cattle lameness detection studies. The table compares these studies based on the number of training and testing frames, the test environment, and the focused activity. As previously mentioned, the dataset consists of a total of 50 videos, including 25 normal and 25 lame cattle videos. Out of these, 30 videos (15 normal and 15 lame) were used for training the model, while the remaining 20 videos (10 normal and 10 lame) were used for testing. Previous studies focused on intensive farming environments, where cattle were primarily managed in indoor grazing conditions and data was collected from specific, limited locations such as within a barn, a passing alley, a narrow path, and corridors related to milking parlors. In contrast, this study focuses on a semi-intensive farming setting and incorporates both indoor and outdoor environments. This approach captures a wider range of cattle movement behaviors, lighting conditions, and surface variations, potentially leading to a more generalizable model for cattle lameness detection.
3.2.1 Data pre-processing
Preprocessing in this investigation was conducted in several distinct steps, including frame extraction and data augmentation. The overall average frame rate was approximately 35.12 frames per second. However, due to the high volume of frames, only 25 images were randomly selected from each training video, resulting in a total of 750 images. These images were then cropped to a 224×224 pixel ratio. However, data augmentation is a widely recognized method for addressing limited data issues and helps to create robust and sustainable classification models. Numerous data augmentation techniques are available, as demonstrated by Jia et al. [4], who employed nine different methods, such as Mixup, Cutout, CutMix, etc.
The primary goal of augmentation in this study is to increase the learning capability of the training model in diverse ways. This study considered only image flipping as an augmentation technique, applied solely to the training data. his decision was driven by several factors, such as:
  • Horizontal flipping realistically simulates left-to-right and right-to-left walking, which are common in cattle locomotion videos. This increases the model’s ability to generalize across different movement orientations without altering the fundamental gait characteristics.
  • Vertical flipping or rotation was also avoided because it inverts the image, resulting in an anatomically impossible scenario with the cattle’s head at the bottom and hooves at the top of the frame. This can hinder the model’s ability to learn realistic gait patterns.
  • The dataset used in this study already includes a variety of backgrounds, as the videos were collected from different online sources. This is why augmentation techniques that artificially alter the background, such as random background replacement, were considered unnecessary and could even introduce misleading variability. Similarly, crop jitter was not utilized due to the risk of removing important contextual information from the video frames.
Therefore, to maintain the integrity of the gait features and avoid introducing misleading transformations, horizontal flipping was considered the most suitable augmentation method for this study. In this continuity, all 30 training videos (750 frames) were horizontally flipped, generating a total of 1500 images. In test data, 500 images were extracted from 20 videos. However, no data augmentation was performed on the test dataset to preserve integrity and ensure a fair assessment of the model’s generalization ability.
3.3 Model Construction
@@
1:
• Training dataset $\mathcal{D}_{\text{train}}$ and testing dataset $\mathcal{D}_{\text{test}}$, where $\mathcal{D}=\{(V_{i},l_{i})\}$,
$V_{i}$ is the $i$-th video, $l_{i}\in\{\text{normal, lame}\}$
• Number of frames per video $N_{\text{frames}}$
• Target size for frame resizing $S_{\text{target}}$
• Data augmentation probability $p_{\text{aug}}$
• Batch size $B$, Number of epochs $E$
2:Trained 3DCNN or ConvLSTM2D model and performance metrics
3:Data Preparation
4:for all $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{test}}$ do
5:  $F_{i}\Leftarrow\text{Pad/Truncate}\big{(}\text{ResizeFrames}(\text{ExtractFrames}(V_{i},\text{max\_frames}=25),S_{\text{target}}),N_{\text{frames}}\big{)}$
6:  if $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}$ then
7:   $(X,y)\Leftarrow(X\cup\{F_{i}\},y\cup\{l_{i}\})$
8:  else
9:   $(X_{\text{test}},y_{\text{test}})\Leftarrow(X_{\text{test}}\cup\{F_{i}\},y_{\text{test}}\cup\{l_{i}\})$
10:  end if
11:end for
12:Apply horizontal flip augmentation to $X_{\text{train}}$ with probability $p_{\text{aug}}$
13:Normalize $X_{\text{train}},X_{\text{test}}$ and one-hot encode labels
14:Model Training
@@
17:Model Evaluation
[…]
20:return $M$, metrics, confusion matrix
@@
Algorithm 1 outlines the pseudocode of the classfication model. It begins with data preparation by processing the training and testing datasets ($\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{test}}$), extracting frames from each video, resizing them to a target size ($S_{\text{target}}$), and ensuring each video has $N_{\text{frames}}$ frames through padding or truncation, creating input data ($X$) and labels ($y$). Training data is augmented with random horizontal flipping (probability $p_{\text{aug}}$). Pixel values are normalized, and behavior labels (“normal” and “lame”) are encoded numerically. Next, model training involves creating a 3D CNN or ConvLSTM2D model, configured using $N_{\text{frames}}$ and $S_{\text{target}}$, and training it on the prepared data ($X_{\text{train}}$, $y_{\text{train}}$) using a batch size ($B$) and epochs ($E$). The model learns patterns indicative of different behavior classes. Finally, model evaluation uses the trained 3D CNN model to predict behavior labels ($\hat{y}$) for the testing dataset ($X_{\text{test}}$), comparing them to true labels ($y_{\text{test}}$) to assess performance. Metrics like accuracy are calculated, and a confusion matrix is generated. The trained model, performance metrics, and confusion matrix are returned as output.
3.3.1 3D CNN